Give the equation of the circle with the given center and radius.
Center at(1,2); radius =1/2
Would I use the formula, y-y^1=m*(x-x1)? Thank you for your help on this.

Answer by nerdybill(5404)

No, your formula is the "point-slope" form of a straight line.
.
The general equation of a circle is:
(x-h)^2 + (y-k)^2 = r^2
where
(h,k) is the center
r is the radius
.
Center at(1,2); radius =1/2
(x-1)^2 + (y-2)^2 = 1/4